How would you calculate this problem in simple interest form made a deposit of $2,000 into an account on August 1 that yields 2% interest compounded annually. How much money will be in that account at the end of 5 years?
Answer by rfer(16322)

A=P(1.02)^5
A=2000(1.02)^5
A=2000*1.104
A=$2208.00